An earthquake generates three kinds of waves: surf
Answer #1

The speed of P waves is approximately 7 km/s, and that of S waves is about 4 km/s

x = ct
=> x = cp*tp = cs*ts

tp = cs/cp * ts

So,

tp - ts = ts(cs/cp - 1)

|tp - ts} = 30 s

= ts((4/7) -1)
= ts(0.43)

30 = ts(0.43)

=> ts = 69.76 s

So,
x = 4 km/s * 69.76
= 280 km (Ans)